OpenStudy (anonymous):

Write equation in slope-intercept form: passes through (-1,4) and (-5,2)

OpenStudy (anonymous):

first, calculate the slope between the two points....

OpenStudy (anonymous):

the slope is -1/2

OpenStudy (wikiemol):

slope is change in y over change in x, so it would be (y1-y2)/(x1-x2). Slope intercept form is y-y1 = m(x-x1) where m is the slope.

OpenStudy (anonymous):

Ok sorry I wasn't at my computer

OpenStudy (anonymous):

the answer is y=1/2x+4.5

OpenStudy (anonymous):

Okay, thanks.

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

Could you help me with another?

OpenStudy (anonymous):

Post it

OpenStudy (anonymous):

slope 2, y-intercept -4

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

y = slope*(x) + y intercept y = 2*x-4

OpenStudy (anonymous):

y=2x-4

OpenStudy (anonymous):

Ok I have two more and thats all: Has an x-intercept of 6 and y-intercept of 3 Passes through (1,2) with slope of -5/3

OpenStudy (anonymous):

y=-5/3x-3

OpenStudy (anonymous):

First, Has an x-intercept of 6 and y-intercept of 3, for that Answer is x+2y-6=0 Second Passes through (1,2) with slope of -5/3, for that Answer is 5x+3y-1 =0

OpenStudy (anonymous):

@ I_suck_at math, once recheck it

OpenStudy (anonymous):

Thank you aswell @SandeepReddy

OpenStudy (anonymous):

Its my pleasure helping out @blah124